OpenStudy (anonymous):

sec(5pi/4) using unit circle/no calc?

OpenStudy (mertsj):

Find the cosine and change it to its reciprocal.

OpenStudy (anonymous):

i know the cos is sqrt2/2 But How would I get the recipricol from 1/sqrt2/2?

OpenStudy (mertsj):

\[\frac{\sqrt{2}}{2}\times \frac{2}{\sqrt{2}}=1\]

OpenStudy (mertsj):

Those numbers are reciprocals because their product is 1. That, by definition, is what reciprocals are: 2 numbers whose product is 1.

OpenStudy (anonymous):

ohhh ok I think I get it now thanks

OpenStudy (mertsj):

yw